Telescope Performance Verification
with IRTC/LFBG - 13-14 May 2009 MST

Observers:: D.Thompson, J.Heidt (mtn)
Telescope Operator: A. Ceranski
Software Support: C. Biddick (phone)
Telescope Support: T. Shih, J. Little

Summary

Tonight was basically a total loss for the following reasons:
  • TCS not delivered in an operable state (OSS left running on lbtmu304 causing conflicts with the centos4 version), 15 min nighttime lost. See IT#2180.
  • M1 was heated to over 13C when it was 10C outside (20C air was being pumped into it!), see IT#2172. This induced poor seeing and poor collimation convergence. 75% of the night lost to this.
  • Incorrect transform was in place and we could not find clear instructions how to fix this. Called Jesper. IT#2181
  • At around 3:52 everything slowed down to the point of unusability. The lbtmu111 server appears to be down. IT#2182

Please have this all fixed by tomorrow night!

Details

TCS not operable at sunset...OSS problems, see IT#. It appears OSS was still running on one of the CentOS5 servers and its connection to the UMAC was causing the CentOS4 OSS to have problems. Lost

/var/log/tcs.log is not readable by LBTO or telescope users, presumably only root? But there are only two messages in there:
May 11 16:52:35 lbtmu107 root: I am debugging
May 11 16:54:16 lbtmu107 root: I am debugging

Wed May 13 19:36:23.228 2009 Preset command started
Wed May 13 19:38:31.193 2009  Preset alert for left IRTC mode ACQUIRE: OK
Wed May 13 19:38:31.244 2009  Preset alert for left IRTC mode ACQUIRE: oss.side[0].secondary_mirror: [moveAbs] UMAC is not ready - brake may be set
Wed May 13 19:38:31.296 2009  Preset command failed for left IRTC mode ACQUIRE from IIF

19:43:34.548 pointto M5_0384 gs=0 Tel="ACQUIRE" PAM- PA=0
The preset succeeded, but the wrong AGW was selected and GCS noted it was not able to get an acquisition image

19:51:45.680 pointto same star, TRACK mode. Star on IRTC.

* Pointing Model Terms ***
IA -314.91330
IE -74.22980
NPAE -12.66810
CA 6.36950
AW -6.65080
AN 22.89390
TF 26.40080
HACA2 -4.10620

IE changed to -70, CA changed to -4

19:56:12.207 pointto BS9148 gs=0 Tel="ACQUIRE" PAM- PA=0 to check. Will collimate here when done. IE changed to -69.5, CA to -6.

Toutside=10.0, T(M1GLASS)=8.5, RH=21% Wind ~5, gusts to 10 m/s.

20:00:23.317 pointto BS9148 gs=0 Tel="ACTIVE". WFS 3096, 1609, ..., 364, 314. FWHM~1.0

Preset stability with time and elevation
20:10:08.121 BS9148 gs= 4 EL=79 031512 00001
20:16:48.168 BS9162 gs= 2 EL=38 032137 00002
20:23:03.865 BS9169 gs=14 EL=29 032727 00003
20:32:07.946 BS9148 gs= 4 EL=83 033741 00004

note In fact, these data will be usable for a collimation model if so desired.

PCS bug found this afternoon that had us using -186.29 instead of the -187.703 value for LUCIFER that we thought we were using. This explains why the IRTC is only 0.3 degrees off rather than the expected ~1.5, PCS restarted to get it to take effect. IE and CA set back to what was determined earlier.

20:43:31.140 BS9148 gs=4 EL=85 034636 00005

Run a dice5 pattern here.
dice5 on BS9148 10x10 30x2sec nloop=5. Field will go through EL~87 about 16 minutes after starting this script. PA=180.18, so source is on AGW_X=0 at radius 188.71 arcsec.
At start: 034831 00006
At end: 042522 00030

21:27:18.947 BS9162 gs= 2 EL=50 043449 00031 FWHM~1.5 arcsec on guider
21:36:12.178 BS9169 gs=14 EL= hit +Z limit on M1, adjust M1 and M2 by -1mm
takeseq here without presetting... 044530 00032
21:47:00.528 BS9169 gs=14 EL=41 044846 00033

Run a dice5 pattern here.
dice5 on BS9169 10x10 30x2sec nloop=10.
At start: 045057 00034
At end: 060443 00083
Two of the five positions are off of the IRTC detector.

23:12:20.490 BS9169 gs=14 EL=53 061541 00084
IE adjusted to -65, CA to 6.8

23:19:53.366 BS9148 gs= 4 EL=61 062453 00085
PA=120 (60 degrees offset from original measurement of 180.18)
Guider seeing 1.4 arcsec FWHM.

dice5 on BS9148 10x10 30x2sec nloop=10 at PA=120
At start: 062813 00086
At end: 074235 00135

23:55 Tout=8.1C T(M1GLASS)=13.2 (SETPOINT problem?!! "integrator" is commanding 20 degrees!) RH=28.1, Wind 6-9m/s. This would explain the lousy seeing. Calling John at 23:38. It was some parameter that was set the previous night. Note added to IT#2172. WFS has been having trouble keeping things collimated (significant fraction of cycles above 400nm).

00:44:02.079 BS9148 gs= 4 EL=45 074608 00136 PA=120

Adjust pointing: IE now -72. CA now -3.0

00:52:26.922 BS9169 gs=14 EL=65 075939 00137 PA=256.
WFS still having trouble, rms typically over 400nm. T(M1GLASS) = 12.0

dice5 on BS9169 10x10 30x2sec nloop=10 at PA=256.
At start: 080130 00138
Aborted at - need to center this field before running the dice5 pattern At end: 080729 00142

The offset in the center script worked fine, but when I did an updatePointingReference command the guide star was lost. This happened twice. See IT# .

dice5 on BS9169 10x8 30x2sec nloop=10 at PA=256. (Yoff=8 arcsec to keep all 5 points on chip)
T(M1GLASS)=11.0, Tout=8.5, guider now seeing 0.8 arcsec FWHM.
At start: 082618 00143
At 02:28 Aaron reports M1 is within 1.5C of ambient. Guider is now seeing 0.7 arcsec FWHM. At end: 094043 00192

02:43:02.303 BS9169 gs=14 EL=67 094525 00193 PA=256.
02:47:28.617 BS9169 gs=14 EL=67 094915 00194 PA=286.
02:43:02.303 BS9169 gs=14 EL=67 095148 00195 PA=316.
02:43:02.303 BS9169 gs=14 EL=67 095416 00196 PA=346.
02:56:30.138 BS9169 gs=14 EL=67 095744 00197 PA=16. (would not take 376)

The on-axis source moves with change of PA using same guide star. There is something not set up correctly in the transform or ??. Checking if this is because of the LEFTZEROPOINT change. Switched back to -187.703 and restarted PCS at 03:04:20. Pointing updated to IE=-82.1, CA=-2.5. Why, after correcting pointing on this source, the subsequent preset is off by ~5 arcsec?

03:07:47.801 BS9169 gs=14 EL=65 101024 00198 PA=16.
03:12:05.873 BS9169 gs=14 EL=65 101418 00199 PA=16.
03:15:50.426 BS9169 gs=14 EL=65 101657 00200 PA=316.
03:22:19.405 BS9169 gs=14 EL=65 102329 00201 PA=286.
03:24:53.830 BS9169 gs=14 EL=65 103013 00202 PA=256.

Neither LEFTZEROPOINT works. The on-axis source moves significantly with changes in PA. We must be using the incorrect transform for this focal station. In going through notes from the recent commissioning logs it is not at all clear what we should be using. 03:39 Called Jesper. We went through the following steps
  • stopAGW -u 0
  • on agw-control as root:
    • /etc/init.d/oacsrv stop
    • /etc/init.d/oacsrv start
  • startAGW -u 0
  • stop GCS
  • start GCS
  • selectAGW AIP_IRTC_L.cfg
  • home the AGW motors

But we did not get a chance to test out if this worked as this was when lbtmu05 started slowing down.

03:57 Twice now (03:52:50.068 and 03:57:08.121) a command issued in iraf for the IRTC hesitated for several tens of seconds before registering on IIFGUI. An ls of /Repository also is taking many tens of seconds to execute from lbtmu05. Again with a preset command from iraf hanging with no response for several tens of seconds. Is this an NFS issue?

04:18:07.274 ACQUIRE - IE now -72, CA now -10.
04:22:27.173 ACTIVE mode. Note pointto command hung again for at least 50 seconds.

04:22:27.173 BS9169 gs=14 EL=57 noimage 00203 PA=316. Single image did not get taken.
This message popped up right when the command continued: "Unknown user: lbcobs". It repeats now for all commands issued in iraf. A shell command (whoami) from within iraf hung for ~65 seconds before continuing. Display has hung now for over 3 minutes. Top shows the load average at 0.01 on lbtmu05. The lbtmu111 server has an orange light on the front panel, and I am unable to ping it. Is this the LDAP or NFS server?

Closing at 04:53. Strong twilight visible on WFS although it and the guider were still working.

-- DavidThompson - 14 May 2009
Topic revision: r8 - 17 May 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback